OpenStudy (anonymous):

Write the equation of the line parallel to 2x + 3y = 1 through points (3,2). Really need to know how to do this

OpenStudy (anonymous):

solve for y and you get \[y= -2/3x +1/3\] Parallel lines have the same slopes.

OpenStudy (anonymous):

use pointe-slope form: y-y_1=m(x-x_1)

OpenStudy (anonymous):

I can write it in y = mx +b form, so y = -2/3x + 1/3 would be correct?

OpenStudy (anonymous):

no. that is the equation of the line 2x+3y=1 plug the point (3,2) into point slope form: y-y_1=m(x-x_1)

OpenStudy (anonymous):

that's y-y divided by 1? what's the underscore?

OpenStudy (anonymous):

Sorry, I've never done this before. Just need to know how to solve one of em and I'll be good to go on the rest. My teacher didn't explain it really at all

OpenStudy (anonymous):

it is y minus the y value of the point (3,2)\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

y-2=-2/3(x-3)

OpenStudy (anonymous):

So you evaluate that?

OpenStudy (anonymous):

yes, solve for y and it should be in the form your teacher wants, unless your teacher wants it in standard form

OpenStudy (anonymous):

So y = -2/3x would be the answer? Just want to make sure I'm on the right track

OpenStudy (anonymous):

no. double check your distribution.

OpenStudy (anonymous):

y-2=-2/3(x-3) y=-2/3x +6/3+2 y=-2/3x+4

OpenStudy (anonymous):

I've gotta get some sleep an am logging off very soon. do you understand it now or have any other questions?

OpenStudy (anonymous):

I think I understand now, thank you very much for your time I appreciate it!

OpenStudy (anonymous):

not a problem. follow me on this and if you ever have trouble again, message me or tag me in a post and I'll help you.

OpenStudy (anonymous):

Thank you! And I should be back every now and then as my class gets more challenging